# Aerial Active STAR-RIS-assisted Satellite-Terrestrial Covert Communications

### v1

@@ V-B7 Practicality Implementation
We conduct some preliminary experiments to demonstrate the feasibility of our approach. Specifically, we utilize a Raspberry Pi 4 Model B, a widely used platform in UAVs and other LAP systems [35], for the experimental implementation, and the specific information is shown in Fig. 11(a). Moreover, Fig. 11(b) presents the inference time of the proposed GDPG algorithm evaluated over different diffusion steps. As can be seen, the inference time linearly increases as the number of denoising steps increases. This trend highlights that the inference time of the proposed GDPG algorithm remains within an acceptable hundred milliseconds. In conclusion, selecting the suitable number of denoising steps depends on finding the appropriate balance between exploration and inference time.

### v2

@@ V-B7 Practicality Implementation
We conduct some preliminary experiments to demonstrate the feasibility of our approach. Specifically, we utilize a Raspberry Pi 4 Model B, a widely used platform in UAVs and other LAP systems [Wilson2022], for the experimental implementation, and the specific information is shown in Fig. 11(a). Moreover, Fig. 11(b) presents the inference time of the proposed GDPG algorithm evaluated over different diffusion steps. As can be seen, the inference time linearly increases as the number of denoising steps increases. With an inference latency under 100 ms, the GDPG algorithm introduces negligible overhead compared to the inherent GEO propagation delay around 250 ms [Chen2020] and the low-frequency control cycles of aerial relay platforms [Real2020], making it suitable for real-time deployment. In conclusion, selecting the suitable number of denoising steps depends on finding the appropriate balance between exploration and inference time.